OpenStudy (anonymous):

On a map of texas the distance between el paso and dallas is 4 inches, the actual distance between these cities is 640 miles, houston and dallas are 1.5 inches apart on the same map how many miles apart houston and dallas

OpenStudy (anonymous):

Dallas-El Paso=4in. (640 mi.) 640/4=160 1 inch=160 mi. 0.5 inch= 160/2=80 mi. 1.5 inch= 160 + 80= 240 mi. 1.5 inch = 240 mi. The distance between Houston and Dallas is 240 miles apart.